OpenStudy (anonymous):

plzzzz help me.. round to the nearest tenth 1.18x + 4.34 = 8.67

OpenStudy (anonymous):

if you help I will reward you:)

OpenStudy (anonymous):

Let's see...

OpenStudy (anonymous):

:) thank you:)

OpenStudy (anonymous):

Okay, so, we have 1.1x + 4.3 = 8.7

OpenStudy (anonymous):

the choices are a.11.0 b.3.8 c.3.7 d.3.6

OpenStudy (anonymous):

Okay, in that case we need to solve, then round it.

OpenStudy (anonymous):

In this case, we have 1.18x + 4.34 = 8.67. Does the question give you anything for x?

OpenStudy (anonymous):

no, we haft to solve for x

OpenStudy (anonymous):

Ah, okay. Then, we would be attempting to get x by itself.

OpenStudy (anonymous):

1.18x + 4.34 = 8.67 Subtract 4.34 from both sides. This gives us... 1.18x = 4.33

OpenStudy (anonymous):

Do you think you can it from there?

OpenStudy (anonymous):

*take it

OpenStudy (anonymous):

yes, and I think you divide 1.18 from 4.33. right?

OpenStudy (anonymous):

Yes sir!

OpenStudy (anonymous):

(Or ma'am.)

OpenStudy (anonymous):

haha im a girl but thts ok:) haha

OpenStudy (anonymous):

Heh. c:

OpenStudy (anonymous):

are you a girl or boy?

OpenStudy (anonymous):

A boy.

OpenStudy (anonymous):

then thank you sir :) lol

OpenStudy (anonymous):

Lol. You're welcome.